# Cogenesis of visible and dark matter in a scotogenic model

### v3

@@ 4 Neutrino mass and LFV constraints
The neutrino mass in the scotogenic model is generated radiatively at one loop as shown in Fig. 2 and explicitly mentioned in Eq. 6. To extract the neutrino oscillation observables one can resort to the Casas-Ibarra parametrization Casas:2001sr for the neutrino mass matrix. Alternatively, we follow the algorithm prescribed in Adhikary:2013bma to find out all the oscillation parameters by constructing the matrix $h=m_{\nu}m_{\nu}^{\dagger}$ which can be diagonalized by a unitary matrix in the following manner
@@
where $m_{1}^{2},m_{2}^{2}$ and $m_{3}^{2}$ are the squared eigenvalues of the neutrino mass matrix. The mixing angles and the CP phase appearing in the unitary matrix can be parametrized using the PDF convention ParticleDataGroup:2024cfk as
@@
where $s_{ij}\equiv\sin\theta_{ij}$ and $c_{ij}\equiv\cos\theta_{ij}$. We have chosen the normal hierarchy i.e. $m_{3}>m_{2}>m_{1}$ to find out the angles and the CP phase while remaining consistent within the $3\sigma$ interval of global fit of the experimental neutrino oscillation data Esteban:2020cvm, cosmological upper bound on sum over neutrino masses i.e $\sum m_{\nu}<0.12$ eV Planck:2018vyg and the stringent constraint on effective neutrino mass parameter from KamLAND-Zen Collaboration i.e $m_{\beta\beta}<(0.061-0.165)$ eV KamLAND-Zen:2016pfg. The benchmark point given in Table 3 satisfies all the neutrino oscillation parameters along with the above mentioned constraints as can be seen from the numerical values quoted in Table 4.
@@
Within the minimal scotogenic model, complex Yukawa couplings can induce processes with lepton flavor violation (LFV) such as $l_{\alpha}\to l_{\beta}\gamma$ and $l_{\alpha}\to 3l_{\beta}$ can occur through one-loop diagrams. Additionally, complex Yukawa couplings can also generate non-zero electric dipole moments (EDMs) of charged leptons via two-loop diagrams. Along with neutrino oscillation constraints, we also implement the experimental constraints arising from such LFV processes. The branching ratio of $l_{\alpha}\to l_{\beta}\gamma$ is given by Toma:2013zsa
@@
The most stringent constraint on such processes arises from the $\mu\to e\gamma$ decay, with the MEG II experiment placing an upper limit on the branching ratio of ${\rm Br}(\mu\to e\gamma)<1.5\times 10^{-13}$ MEGII:2025gzr. For $l_{\alpha}\to 3l_{\beta}$ processes, both the 1-loop box and penguin diagrams can contribute, however, we have not considered penguin diagrams involving the Higgs boson, since the corresponding Yukawa couplings are suppressed. Although this suppression does not apply to processes involving $\tau$-leptons, the experimental limits on LFV in the $\tau$ sector are comparatively weaker, making this approximation valid for our analysis. The branching ratio for the $l_{\alpha}\to 3l_{\beta}$ process can be expressed as Toma:2013zsa
@@
and the form of the coefficient $B$ is taken from Toma:2013zsa. The strongest constraint on these decays comes from the process $\mu\to 3e$, with the current limit from the SINDRUM experiment being ${\rm Br}(\mu\to 3e)<10^{-12}$ SINDRUM:1987nra. The Mu3e experiment is expected to improve this sensitivity to ${\rm Br}(\mu\to 3e)<10^{-16}$ in the future Amarinei:2025ntv. In our analysis, we have required the branching ratio to remain below the projected Mu3e sensitivity, which automatically satisfies the present experimental bound.
Along with the LFV processes, the CP violating Yukawa couplings present in our framework can introduce EDMs to charged leptons, which can be written as Abada:2018zra
@@
where the Majorana and Dirac quartic invariants are denoted respectively as $J^{M}_{ij\alpha\beta}=\mathrm{Im}\left[y_{j\alpha}^{*}\,y_{j\beta}^{*}\,y_{i\beta}\,y_{i\alpha}\right]$ and $J^{D}_{ij\alpha\beta}=\mathrm{Im}\left[y_{j\alpha}^{*}\,y_{j\beta}\,y_{i\beta}^{*}\,y_{i\alpha}\right]$ with the 1-loop functions denoted by $I_{M}(\xi_{i},\xi_{j})$ and $I_{D}(\xi_{i},\xi_{j})$. The upper limit on the electron EDM is at least ten orders of magnitude more stringent than for other charged leptons, with the latest bound given by the ACME Collaboration as $\left|d_{e}\right|<1.1\times 10^{-29}\,e\,{\rm cm}$ ACME:2018yjb.
@@
Apart from these laboratory limits, there are constraints on the total mass of all the neutrino species ($\sum m_{\nu}$) from Planck Planck:2018vyg as can be seen from Fig. 7 where the allowed points are shown in colour. However, satisfying all the neutrino oscillation parameters along with leptogenesis and relic density while remaining consistent with the structure formation bounds significantly shrinks the allowed parameter region as is shown in Fig. 8. Relaxing the mass scales of the heavy dark sector particles might open up new regions of parameter space but this leads to nothing new insights on the cogenesis framework for which we prefer to fix the mass scales.
Figure 8: The gray dots represent the allowed parameter space satisfying correct baryon asymmetry and DM relic abundance while obeying the FIMP lower mass bound. The coloured points additionally satisfies all the neutrino oscillation observables, cosmological upper bound on $\sum m_{\nu}$ and LFV constraints. The benchmark point given in Table 3 is indicated by ✖.
@@ Acknowledgements.
We thank Deep Ghosh and Tarak Nath Maity for useful discussions. TSR and RP acknowledges the Department of Science and Technology, Government of India, under the Grant Agreement No. MTR/2023/000469 (MATRICS) for financial assistance.

### v4

@@ 4 Neutrino mass and LFV constraints
The neutrino mass in the scotogenic model is generated radiatively at one loop as shown in Fig. 2 and explicitly mentioned in Eq. 6. To extract the neutrino oscillation observables one can resort to the Casas-Ibarra parametrization Casas and Ibarra (2001) for the neutrino mass matrix. Alternatively, we follow the algorithm prescribed in Adhikary et al. (2013) to find out all the oscillation parameters by constructing the matrix $h=m_{\nu}m_{\nu}^{\dagger}$ which can be diagonalized by a unitary matrix in the following manner
@@
where $m_{1}^{2},m_{2}^{2}$ and $m_{3}^{2}$ are the squared eigenvalues of the neutrino mass matrix. The mixing angles and the CP phase appearing in the unitary matrix can be parametrized using the PDF convention Navas and others (2024) as
@@
where $s_{ij}\equiv\sin\theta_{ij}$ and $c_{ij}\equiv\cos\theta_{ij}$. We have chosen the normal hierarchy i.e. $m_{3}>m_{2}>m_{1}$ to find out the angles and the CP phase while remaining consistent within the $3\sigma$ interval of global fit of the experimental neutrino oscillation data Esteban et al. (2020), cosmological upper bound on sum over neutrino masses i.e $\sum m_{\nu}<0.12$ eV Aghanim and others (2020) and the stringent constraint on effective neutrino mass parameter from KamLAND-Zen Collaboration i.e $m_{\beta\beta}<(0.061-0.165)$ eV Gando and others (2016). The benchmark point given in Table 3 satisfies all the neutrino oscillation parameters along with the above mentioned constraints as can be seen from the numerical values quoted in Table 4.
@@
Within the minimal scotogenic model, complex Yukawa couplings can induce processes with lepton flavor violation (LFV) such as $l_{\alpha}\to l_{\beta}\gamma$ and $l_{\alpha}\to 3l_{\beta}$ can occur through one-loop diagrams. Additionally, complex Yukawa couplings can also generate non-zero electric dipole moments (EDMs) of charged leptons via two-loop diagrams. Along with neutrino oscillation constraints, we also implement the experimental constraints arising from such LFV processes. The branching ratio of $l_{\alpha}\to l_{\beta}\gamma$ is given by Toma and Vicente (2014)
@@
The most stringent constraint on such processes arises from the $\mu\to e\gamma$ decay, with the MEG II experiment placing an upper limit on the branching ratio of ${\rm Br}(\mu\to e\gamma)<1.5\times 10^{-13}$ Afanaciev and others (2025). For $l_{\alpha}\to 3l_{\beta}$ processes, both the 1-loop box and penguin diagrams can contribute, however, we have not considered penguin diagrams involving the Higgs boson, since the corresponding Yukawa couplings are suppressed. Although this suppression does not apply to processes involving $\tau$-leptons, the experimental limits on LFV in the $\tau$ sector are comparatively weaker, making this approximation valid for our analysis. The branching ratio for the $l_{\alpha}\to 3l_{\beta}$ process can be expressed as Toma and Vicente (2014)
@@
and the form of the coefficient $B$ is taken from Toma and Vicente (2014). The strongest constraint on these decays comes from the process $\mu\to 3e$, with the current limit from the SINDRUM experiment being ${\rm Br}(\mu\to 3e)<10^{-12}$ Bellgardt and others (1988). The Mu3e experiment is expected to improve this sensitivity to ${\rm Br}(\mu\to 3e)<10^{-16}$ in the future Amarinei (2025). In our analysis, we have required the branching ratio to remain below the projected Mu3e sensitivity, which automatically satisfies the present experimental bound.
Along with the LFV processes, the CP violating Yukawa couplings present in our framework can introduce EDMs to charged leptons, which can be written as Abada and Toma (2018)
@@
where the Majorana and Dirac quartic invariants are denoted respectively as $J^{M}_{ij\alpha\beta}=\mathrm{Im}\left[Y_{j\alpha}^{*}\,Y_{j\beta}^{*}\,Y_{i\beta}\,Y_{i\alpha}\right]$ and $J^{D}_{ij\alpha\beta}=\mathrm{Im}\left[Y_{j\alpha}^{*}\,Y_{j\beta}\,Y_{i\beta}^{*}\,Y_{i\alpha}\right]$ with the 1-loop functions denoted by $I_{M}(\xi_{i},\xi_{j})$ and $I_{D}(\xi_{i},\xi_{j})$. The upper limit on the electron EDM is at least ten orders of magnitude more stringent than for other charged leptons, with the latest bound given by the ACME Collaboration as $\left|d_{e}\right|<1.1\times 10^{-29}\,e\,{\rm cm}$ Andreev and others (2018).
@@
Figure 8: Gray dots represent the allowed parameter space satisfying correct baryon asymmetry and DM relic abundance while obeying the FIMP lower mass bound presented in terms of the Yukawa couplings of (8(a)) $N_{1}$, (8(b)) $N_{2}$, (8(c)) $N_{3}$ RHN and (8(d)) quartic couplings between Higgs and the inert scalar doublet. The coloured points additionally satisfies all the neutrino oscillation observables, cosmological upper bound on $\sum m_{\nu}$ and LFV constraints. The benchmark point given in Table 3 is indicated by ✖.
Apart from these laboratory limits, there are constraints on the total mass of all the neutrino species ($\sum m_{\nu}$) from Planck Aghanim and others (2020) as can be seen from Fig. 7 where the allowed points are shown in colour. However, satisfying all the neutrino oscillation parameters along with leptogenesis and relic density while remaining consistent with the structure formation bounds significantly shrinks the allowed parameter region as is shown in Fig. 8. Additionally, it is evident from the appearance of gray points in Fig. 8(d) that the parameter region with $\lambda_{3}\lesssim 3$ is highly constrained due to the stringent structure formation bound on the mass of freeze-in DM. Relaxing the mass scales of the heavy dark sector particles might open up new regions of parameter space but this leads to nothing new insights on the cogenesis framework for which we prefer to fix the mass scales.
@@ Acknowledgements.
We thank Deep Ghosh and Tarak Nath Maity for useful discussions. T. S. R. and R. P. acknowledge the Department of Science and Technology, Government of India, under the ANRF Grant Agreement No. MTR/2023/000469 (MATRICS) for financial assistance.
Appendix A Benchmark point with higher value of $\lambda_{3}$
Yukawa couplings ($Y_{\alpha i}$)
[…]
Table 5: Details of the benchmark point for the cogenesis scenario with higher value of $\lambda_{3}$.
(a) Same as Fig. 5.
(b) Same as Fig. 4(a).
Figure 9:
In this appendix, we present an additional benchmark point with higher values of $\lambda_{3}$. The mass scales of the heavy neutrinos and the inert doublet are kept same as the benchmark presented in Table 3 with $\lambda_{3}=8.114$, $\lambda_{5}=3.046\times 10^{-2}$ and the Yukawa couplings shown in Table 5. The evolution of various abundances and the relevant reaction densities are given in Figs. 9(a) and 9(b) respectively. Due to the increased thermal scatterings with the primordial bath, the decay of the inert scalar doublet is delayed resulting in a higher DM mass of $M_{N_{1}}=9.6$ MeV, while remaining consistent with the structure formation bounds along with other constraints discussed in the main text.